# Revisit, Extend, and Enhance Hessian-Free Influence Functions

### v2

@@ 4 Correctness Verification on Synthetic Data
5 Noisy Label Correction for Vision Datasets
In this section, we demonstrate the effectiveness of our IP Ensemble in identifying detrimental samples on noisy vision datasets. Specifically, we choose three benchmark datasets CIFAR-10N [60], CIFAR-100N [60], and Animal-10N datasets [50] in the noisy label learning area. CIFAR-10N encompasses three distinct noise settings: aggregate, random, and worst, denoted as “-a,” “-r,” and “-w,” respectively. “a” means that labels are derived via majority voting among three annotators, with ties being resolved randomly, “r” adopts the label provided by the first annotator, while “w” selects the label from the least reliable annotator.
For competitive methods, we choose the following influence function-based methods. TracIn [46] replaces the Hessian matrix with the identity matrix and considers checkpoints during the training process; LiSSA [26] and EKFAC [19] employ implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [31] swaps the order of the matrix multiple for obtaining a closed-form expression; Self-TracIn [52] and Self-LiSSA [5] are the self-expression versions of TracIn and LiSSA, where is replaced with and only the last checkpoint, i.e., the converged model parameters, is used in Self-TracIn. GEX [25] utilizes ensemble methods based on checkpoints from extra stochastic gradient descent on the converged model. TDA [3] focuses checkpoints during the training process, ensembles the influence via EKFAC [19]. Our IP Ensemble method constitutes an ensemble version of IP with dropout applied on model parameters and an ensemble size of 5.
@@ 5 Noisy Label Correction for Vision Datasets
To reduce randomness, we train a ResNet-34 network [22] once on each dataset to establish a baseline model. Subsequently, based on the same model, we employ the aforementioned influence function-based methods to identify 5% of detrimental samples. Following this, we conduct five retraining iterations of the ResNet-34 network, each time removing the identified detrimental samples from the training set. We report the average accuracy and standard deviation of the above influence function-based methods across these five retrainings in Table 1. In general, these influence function-based methods are effective in identifying detrimental samples. Upon retraining the ResNet-34 model without these identified samples, nearly every result obtained by these methods outperforms the vanilla ResNet-34 trained on the entire dataset, except for EKFAC on CIFAR-10N-w. Notably, our IP Ensemble consistently outperforms other baseline methods across various noise conditions and datasets. Particularly noteworthy is the performance of our IP Ensemble in the most challenging scenario, CIFAR-100N, achieving the highest recorded accuracy of 62.25% on the test set, surpassing the vanilla cross-entropy accuracy of 56.41%. Moreover, TDA and TracIn only achieve moderate performance because they introduce multiple checkpoints throughout the training process. These approaches might dilute the effectiveness of identifying truly detrimental samples for the converged model, as the influence of each checkpoint may vary. Additionally, compared to IP, IP Ensemble delivers further performance gains, highlighting the benefits of the ensemble strategy for enhancing model generalization. Moreover, the average accuracy of the IP Ensemble on the test set reaches its peak at 82.93%, surpassing both the vanilla cross-entropy performance of 80.90% and the second-best accuracy of Self-TracIn at 82.45%. More experimental results and analysis on different percentages of removed samples, parameter analysis on different dropout rates, ensemble sizes, and base model architectures can be found in Appendix C.2- C.4, respectively.
@@
In addition, we also present the running times of these influence function-based methods. Despite some baselines having linear time complexity, there is significant divergence in their real execution times. Given that our IP exhibits exceptional speed and similar execution times, we consider them as the baseline and compute the speed improvement factors over other baseline methods, as depicted in Figure 4. For ensemble methods including TracIN, TDA, GEX, and our IP Ensemble, parallel computation can be applied to accelerate the running time if enough resources are allowed; if calculated serially, the time grows linearly with the ensemble size. We do not report their running time in Figure 4. With the exception of Self-TracIn, our IP runs over 100 times faster than LiSSA, EKFAC, DataInf, and Self-LiSSA. Notably, on Animal-10N, IP is over 800 times faster than EKFAC. It is worth noting that Self-TracIn and Self-LiSSA are slower than their standard versions due to the fixed when calculating the influence of each sample. The simplicity and efficiency of IP make it a promising tool for analyzing sample influence on deep models. The time complexities and execution time of these methods can be found in Appendix C.5.
@@ 6 Data Curation Towards Fine-Tuning of Fairer NLP Models
In this experiment, we employ three datasets—RTE, COLA, and QLIN—from the GLUE repository [56] to fine-tune the widely-used language model RoBERTa [35]. Our focus is on group fairness, necessitating that machine learning models treat samples within various predefined subgroups comparably. To assess fairness, we adopt the methodology outlined in [47], which involves perturbing the demographic information within each sample and scrutinizing whether the model yields identical predictions for the original sample and its corresponding perturbed counterpart . This evaluation metric "fair score" is defined as over all the samples in the test set, normalized by the size of the test set, where is the model predictor. Note that the fair score is a negative metric, so smaller values are preferable. Utility and fairness serve as distinct perspectives for evaluating the performance of a model. Thus, in our fine-tuning experiments, we consider both utility and fairness. Employing the same influence function-based methods and our IP Ensemble as in the previous section, we conduct a comparative analysis. For each method, we calculate the influence on both utility and fairness within the fine-tuning set. Subsequently, We identify and remove the 5% most detrimental samples in terms of utility and fairness before fine-tuning the RoBERTa model to optimize performance across both accuracy and fairness metrics.
@@
The results of this experiment are presented in Figure 5, where the Y-axis for fairness is inverted. The brown crossing denotes the performance of using all the samples for model fine-tuning as the base model. Building upon this, we plot a horizontal and a vertical line in each figure and divide the space by fairness and utility results into four regions. The green area in the top right corner signifies a model that is both fairer and more accurate compared to the baseline model. Results within this green region can be considered Pareto improvements, enhancing both utility and fairness simultaneously. It is evident most results are located in the green area, indicating the existence of detrimental samples, and not all the samples are helpful to the model performance. This also implies that influence function methods are effective to identifying the detrimental samples, even for non-convex deep models. However, it is important to note that some competitive methods yield a trade-off or even Pareto deterioration. For instance, LiSSA demonstrates a better fair score but worse accuracy compared to the base model on RTE; conversely, it exhibits better accuracy but worse fairness on QLIN. EKFAC shows similar performance on COLA and QLIN. Self-LiSSA demonstrates Pareto deterioration on COLA. We hypothesize that the approximation of the inverse of the Hessian matrix may suffer from large errors, leading to heavily divergent influence estimations from the true values. Our IP Ensemble consistently achieves Pareto improvements across all three datasets, yielding the best results compared to other influence-based methods.
7 Defense Against Adaptive Evasion Adversaries
@@
In this experiment, we utilize a Logistic Regression model with three datasets: Bank [40], CelebA [37], and JigsawToxicity [41]. Following the protocol in Section 3, we consider a white-box adversary [39] to craft adversarial samples. For each sample in the validation set , we perturb it by changing the feature and keeping unchanged. The attacker perturbs between 5% and 25% of the test set samples at random. By quantifying the impact of samples on model robustness, we trim 5% detrimental samples in the training set through influence functions. The boxplot depicted in Figure 6 demonstrates the performance variations across various influence function-based methods over 10 distinct attacks. Since the Logistic Regression model is convex, TDA simplifies to EKFAC, and TracIn and GEX simplify to IP; therefore, we will omit their duplicate results in this figure and the following table. In general, influence function-based methods are highly effective against adaptive evasion attackers, showcasing superior resilience particularly post-attack, notably evident in scenarios involving Bank and JigsawToxicity. Among these influence-based methods, our IP Ensemble demonstrates competitive efficacy with the best or the second best on all the datasets.
@@ 8 Conclusion
In this paper, we revisited, extended, and enhanced the TracIn method into the Inner Product (IP) formulation. By proving the exhibition of order-consistency in several scenarios between IP and influence function, we demonstrated substituting the inverse of the Hessian matrix with an identity matrix offers a practical and computationally efficient solution to estimating sample influence. Based on that, we extended our IP to measure the sample influence on fairness and robustness. Continually, we enhanced the generalization of IP by introducing an ensemble strategy. We verified the correctness of IP on synthetic datasets and extensive evaluations on noisy label detection, data curation for fair NLP model fine-tuning, and defense against adaptive adversarial attacks. Overall, our IP Ensemble highlighted the potential of simple, yet powerful, approximations in influence estimation.
@@
Appendix A Rationality of Inner Product
In Figure 1, we visualize the direction of and , with the influence scores roughly representing the angle between the sample gradient and or , where induces a rotation of between these two directions. For a complete analysis, we further split Region II and IV into Sub-Regions as shown in Figure 7 assuming . Given gradients of two samples and , we analyze all possible scenarios as follows.
(I, I). In the scenario where both and have gradients in Region I, implies , which is the opposite of order-consistency. While these samples are jointly recognized by and as detrimental, there is little practical difference.
(I, II). In the scenario where has gradients in Region I and has gradients in Region II, the order-consistency partially holds, depending on which Sub-Region the gradients of is in. For Sub-Region and , we have and , and the order-consistency holds. For Sub-Region , however, and , thus the order-consistency partially holds. While samples in Sub-Region are jointly recognized by and as detrimental, there is little practical difference.
(I, III). In the scenario where has gradients in Region I and has gradients in Region III, and , therefore the order-consistency holds.
(II, IV). In the scenario where has gradients in Region II and displays gradients in Region IV, the concept of order-consistency is partially maintained.
(II, II). In the scenario where both and have gradients in Region II, implies , the order-consistency holds.
Based on the symmetry in samples ( and ) and regions (I and III, II and IV), we can get all order-consistencies in Table 3. The bold parts are discussed above, and the rest parts are from the symmetry. In all scenarios where order-consistency does not hold, samples are jointly recognized by both and as either beneficial or detrimental, resulting in minimal practical difference.
| Gradient of | Gradient of | Order-Consistency |
| Region I | Region I | Not Hold |
| Region I | Region II, IV | Partially Hold |
| Region I | Region III | Hold |
| Region II | Region I, III, IV | Partially Hold |
| Region II | Region II | Hold |
| Region III | Region III | Not Hold |
| Region III | Region II, IV | Partially Hold |
| Region III | Region I | Hold |
| Region IV | Region I, II, III | Partially Hold |
| Region IV | Region IV | Hold |
Appendix B Detailed Information on Datasets and Model Training
We describe dataset details, model training, and other information used in the main paper, below.
B.1 Datasets
We cover our generated synthetic datasets in Section 4, vision datasets with noisy labels in Section 5, text datasets in Section 6, and datasets for robustness in Section 7.
B.1.1 Synthetic Datasets
We generate two synthetic datasets to validate the correctness of our IP method in measuring the convex and non-convex models’ utility, fairness, and robustness. Specifically, we generate a linear dataset using the scikit-learn [45] library’s make_blobs function, which consists of 150 training samples and 100 test samples. The second dataset is the non-linear half moons dataset so that we can train a Multi-Layer Perception network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. Here we manually flip the labels of 10 samples from each class to add noise to the data.
B.1.2 Vision Datasets
Both the CIFAR-10N and CIFAR-100N datasets [43] consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets [29], respectively. Each input is a 32x32 RGB image with a dimension of (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50,000 image-label pairs, and the test set contains 10,000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. Due to the substantial number of labeling classes in CIFAR-100N, only a single noisy setting is available, with an overall noise rate of 40.20%.
B.1.3 NLP Datasets
For the three GLUE datasets—RTE, CoLA, and QNLI[56], we perturbe the validation set as follows. The Recognizing Textual Entailment (RTE) dataset consists of sentence pairs labeled as entailment or not entailment. It is derived from a series of annual textual entailment challenges. The dataset includes 2,490 training examples and 277 validation examples. Since the test set does not have labels, we split the validation set into two parts, with one half used as the validation set for computing influence, and the other half used as the test set. Subsequently, we generated perturbed datasets for both the validation set and the test set using a seq2seq model [47]. The Corpus of Linguistic Acceptability (CoLA) dataset consists of sentences labeled as grammatically acceptable or unacceptable. This dataset is derived from publications in linguistics and includes 8,551 training examples and 1,043 validation examples. As with RTE, the test set for CoLA does not have labels, so we split the validation set into two parts, using one half as the validation set for computing influence, and the other half as the test set. Perturbed datasets were generated using a seq2seq model [47] for both the validation set and the test set. The Question Natural Language Inference (QNLI) dataset is a large-scale corpus for question answering, consisting of question-sentence pairs from the Stanford Question Answering dataset. The task is to determine whether the context sentence contains the answer to the question. It includes 104,743 training examples and 5,463 validation examples. Similar to the other datasets, the test set does not have labels, so we split the validation set into two parts, using one half as the validation set for computing influence, and the other half as the test set. Perturbed datasets were also generated using a seq2seq model [47] for both the validation set and the test set. These datasets represent diverse natural language understanding tasks, and using them helps evaluate the fairness and utility of the fine-tuned RoBERTa model.
B.1.4 Datasets for Defencing Attacks
We utilize three datasets: Bank, CelebA, and JigsawToxicity, to evaluate the defense against adaptive evasion adversaries. The Bank dataset [40] consists of features extracted from direct marketing campaigns of a Portuguese banking institution. The goal is to predict whether a client will subscribe to a term deposit. The dataset includes 18,292 training examples, 6,098 validation examples, and 6,098 test examples. The feature dimension of the dataset is 51. Both the validation and test sets have labels. The CelebA dataset [37] is a large-scale face attributes dataset with more than 200,000 celebrity images, each annotated with 40 binary attributes. For this experiment, we focus on a subset of these images. The dataset is split into 62,497 training examples, 20,833 validation examples, and 20,833 test examples. The feature dimension of the dataset is 39. Both the validation and test sets have labels. The JigsawToxicity dataset [41] contains a large number of comments from Wikipedia labeled for toxicity. The goal is to predict the toxicity level of a given comment. The dataset includes 18,000 training examples, 6,000 validation examples, and 6,000 test examples. The feature dimension of the dataset is 385. These datasets cover a range of tasks from marketing prediction and face attribute recognition to toxicity detection, providing a robust evaluation of the model’s defense mechanisms against adaptive adversaries.
B.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 [22] architecture used as the base model for the noisy vision datasets, then the RoBERTa [35] NLP transformer model. We also describe implementation details and parameter values for the label correction baselines used in Sections 7 and the influence-based baselines used throughout the paper.
B.2.1 MLP
In our experiments as described in Section 4, we utilized a three-layer Multi-Layer Perceptron (MLP) classifier. This MLP is structured with three fully connected layers, each followed by a ReLU activation function. Specifically, the input layer maps the input features to 32 neurons, the hidden layer maintains this dimensionality, and the output layer is a single neuron producing the final prediction. The final layer’s output is passed through a sigmoid activation function, generating a probability score for prediction.
B.2.2 ResNet-34
The ResNet-34 model in Section 5 was proposed in [22] and is a 34-layer convolutional neural network pretrained on the ImageNet-1K dataset at resolution 224 224. The pretrained model block is fine-tuned on the CIFAR-10N/CIFAR-100N training set experiments with default parameters– minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), and number of epochs (100), for all experiments.
B.2.3 RoBERTa
We use the Roberta-base model from Huggingface222https://huggingface.co/docs/transformers/model_doc/roberta in Section 6. The learning rate is 0.00001 and the batch sizes of RTE, CoLA, and QNLI are 64, 16, and 32. The model is fine-tuned over 10 epochs. The loss function used is a negative log-likelihood as the datasets are all for binary classification.
B.2.4 Logistic Regression
B.2.5 Influence-Based Baselines
In our experiments, we utilize the following influence function-based methods as baselines: TracIn [46] replaces the Hessian matrix with the identity matrix; LiSSA [26] and EKFAC [19] use implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [31] swaps the order of matrix multiplication to obtain a closed-form expression; Self-TracIn [52] and Self-LiSSA [5] are self-expression versions of TracIn and LiSSA, where is replaced with . GEX [25] utilizes ensemble methods based on checkpoints from extra stochastic gradient descent on the converged model. TDA [3] focuses checkpoints during the training process, ensembles the influence via EKFAC [19]. Our IP closely resembles TracIn and GEX, with the added application of dropout with on the model parameters. TracIn, GEX, IP, and IP Ensemble are all Hessian-free and they are ensemble versions, except for IP. Here, we present their calculations to highlight the differences between them as follows:
where in IP is the converged model parameters; denotes the set of model parameters from the saved checkpoints during model training; is obtained by training the converged model for several extra iterations; IP Ensemble gets by applying dropout mechanisms on the model; is the ensemble size. The advantages of IP Ensemble are that it does not require recording checkpoints compared to TracIn, and it does not need extra model training processes compared to GEX. We argue that the early checkpoints might not be helpful to analyze the sample influence and the extra iterations on the converged model make little effects.
Appendix C Additional Experimental Results and Analysis
C.1 IP for measuring fairness and robustness

### v4

@@ 5 Noisy Label Correction for Vision Datasets
To reduce randomness, we train a ResNet-34 network [23], which is a 34-layer convolutional neural network pretrained on the ImageNet-1K dataset at resolution 224 224. The pretrained model block is fine-tuned on the CIFAR-10N/CIFAR-100N training set using default parameters: minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), and number of epochs (100). Subsequently, based on the same model, we employ the aforementioned influence function-based methods to identify 5% of detrimental samples. Following this, we conduct five retraining iterations of the ResNet-34 network, each time removing the identified detrimental samples from the training set. For each method, we repeat the following procedure five times (each time with an independent run): we first identify the bottom 5% most detrimental samples, then retrain the model on the remaining 95% of the training data. Tables 1 and 5 report the results averaged over these five runs. Table 1 reports the average accuracy and standard deviation of the above influence function-based methods across these five retrainings. In general, these influence function-based methods are effective in identifying detrimental samples at different levels. Upon retraining the ResNet-34 model without these identified samples, nearly every result obtained by these methods outperforms the vanilla ResNet-34 trained on the entire dataset, except for EKFAC on CIFAR-10N-w. Moreover, TracIn and TDA achieve only moderate performance, largely due to their reliance on multiple checkpoints throughout the training process. This highlights why we believe that simply summing up sample influence across different checkpoints fails to effectively capture the dynamic and evolving value of a sample during training.
@@
Notably, our IP Ensemble consistently outperforms other baseline methods across various noise conditions and datasets. Particularly noteworthy is the performance of our IP Ensemble in the most challenging scenario, CIFAR-100N, achieving the highest recorded accuracy of 62.25% on the test set, surpassing the vanilla cross-entropy accuracy of 56.41%. Compared to IP, IP Ensemble delivers further performance gains, highlighting the benefits of the ensemble strategy for enhancing model generalization. Moreover, the average accuracy of the IP Ensemble on the test set reaches its peak at 82.93%, surpassing both the vanilla cross-entropy performance of 80.90% and the second-best accuracy of Self-TracIn at 82.45%. More experimental results and analysis on different percentages of removed samples, parameter analysis on different dropout rates, ensemble sizes, and base model architectures can be found in Appendix A.2- A.4.
@@
In addition, we also present the running times of these influence function-based methods. Despite some baselines having linear time complexity, there is significant divergence in their real execution times. Given that our IP exhibits exceptional speed and similar execution times, we consider them as the baseline and compute the speed improvement factors over other baseline methods, as depicted in Figure 2. For ensemble methods including TracIN, TDA, GEX, and our IP Ensemble, parallel computation can be applied to accelerate the running time if enough resources are allowed; if calculated serially, the time grows linearly with the ensemble size. When we visualize their running time in the logarithm scale, the extended time due to the ensemble size is not significant; for example, EKFAC and its ensemble version TDA. Thus, we do not report the ensemble-based methods in Figure 2, except for the above example of TDA. With the exception of Self-TracIn, our IP runs over 100 times faster than LiSSA, EKFAC, DataInf, and Self-LiSSA. Notably, on Animal-10N, IP is over 800 times faster than EKFAC. The time complexities and execution time of these methods can be found in Appendix A.5.
@@ 6 Data Curation Towards Fine-Tuning of Fairer NLP Models
In this experiment, we employ three datasets—RTE, CoLA, and QNLI—from the GLUE repository [57] to fine-tune the widely-used language model RoBERTa [36]. Our focus is on group fairness, necessitating that machine learning models treat samples within various predefined subgroups comparably. These datasets represent diverse natural language understanding tasks, thus allowing us to comprehensively evaluate the fairness and utility of the fine-tuned RoBERTa model. The Recognizing Textual Entailment (RTE) dataset consists of sentence pairs labeled as entailment or non-entailment, derived from a series of textual entailment challenges, containing 2,490 training examples and 277 validation examples. The Corpus of Linguistic Acceptability (CoLA) dataset includes sentences labeled as grammatically acceptable or unacceptable, derived from linguistic publications, and consists of 8,551 training examples and 1,043 validation examples. The Question Natural Language Inference (QNLI) dataset is a large-scale corpus for question answering, comprising question-sentence pairs from the Stanford Question Answering dataset, with 104,743 training examples and 5,463 validation examples. As the test sets for these datasets lack labels, we split each validation set into two equal parts, utilizing one half as the validation set for computing influence, and the other half as the test set. Perturbed versions of both the validation and test sets are generated using a seq2seq model as detailed in Qian et al. [48]. To assess fairness, we adopt the methodology outlined in Qian et al. [48], which involves perturbing the demographic information within each sample and scrutinizing whether the model yields identical predictions for the original sample and its corresponding perturbed counterpart . The fairness evaluation metric ("fair score") is defined as over all test samples, normalized by the test set size, where is the model predictor. Note that the fair score is a negative metric; hence, smaller values are preferable. We fine-tune the RoBERTa-base model from Huggingface111https://huggingface.co/docs/transformers/model_doc/roberta with the following experimental settings: a learning rate of , batch sizes of 64, 16, and 32 for RTE, CoLA, and QNLI, respectively, and a total of 10 epochs. The loss function used is negative log-likelihood, appropriate for these binary classification tasks.
Utility and fairness serve as distinct perspectives for evaluating model performance. Thus, in our fine-tuning experiments, we consider both dimensions. Employing the same influence function-based methods and our IP Ensemble as in the previous section, we conduct comparative analyses. For each method, we compute influence scores for both utility and fairness on the fine-tuning set, aggregate them into a single joint ranking, and remove the bottom 5% samples under this joint criterion before fine-tuning the RoBERTa model.
@@
The results of this experiment are presented in Figure 3, where the Y-axis for fairness is inverted. The brown crossing denotes the performance of using all the samples for model fine-tuning as the base model. Building upon this, we plot a horizontal and a vertical line in each figure and divide the space by fairness and utility results into four regions. The green area in the top right corner signifies a model that is both fairer and more accurate compared to the baseline model. Results within this green region can be considered Pareto improvements, enhancing both utility and fairness simultaneously. It is evident most results are located in the green area, indicating the existence of detrimental samples, and not all the samples are helpful to the model performance. This also implies that influence function methods are effective to identifying the detrimental samples, even for non-convex deep models. However, some competitive methods yield a trade-off or even Pareto deterioration. For instance, LiSSA demonstrates a better fair score but worse accuracy compared to the base model on RTE; conversely, it exhibits better accuracy but worse fairness on QLIN. EKFAC shows similar performance on COLA and QLIN. Self-LiSSA demonstrates Pareto deterioration on COLA. Our IP Ensemble consistently achieves Pareto improvements across all three datasets, often yielding the best results compared to other methods.
7 Defense Against Adaptive Adversaries
@@
In this experiment, we utilize a Logistic Regression model with three datasets: Bank [41], CelebA [38], and JigsawToxicity [42] to evaluate the defense against adaptive evasion adversaries. The Logistic Regression model is implemented using the sklearn library. The model uses L2 regularization with and a maximum iteration limit of 2,048 [11]. The Bank dataset consists of features extracted from direct marketing campaigns of a Portuguese banking institution, aiming to predict whether a client will subscribe to a term deposit, including 18,292 training examples, 6,098 validation examples, and 6,098 test examples, with a feature dimension of 51. The CelebA dataset is a large-scale face attributes dataset with more than 200,000 celebrity images annotated with 40 binary attributes; we focus on a subset comprising 62,497 training examples, 20,833 validation examples, and 20,833 test examples, each with a feature dimension of 39. The JigsawToxicity dataset includes comments from Wikipedia labeled for toxicity prediction, containing 18,000 training examples, 6,000 validation examples, and 6,000 test examples with a feature dimension of 385. Following the protocol in Section 3, we consider a white-box adversary [40] to craft adversarial samples. For each sample in the validation set , we perturb it by changing the feature and keeping unchanged. The attacker perturbs between 5% and 25% of the test set samples at random. By quantifying the impact of samples on model robustness, we trim 5% detrimental samples in the training set through influence functions. The boxplot depicted in Figure 4 demonstrates the performance variations across various influence function-based methods over 10 distinct attacks. Since the Logistic Regression model is convex, TDA simplifies to EKFAC, and TracIn and GEX simplify to IP; therefore, we omit their duplicate results in this figure and the following table. For the convex case, the existence of the inverse Hessian matrix is guaranteed, and LiSSA is the most suitable algorithm for influence estimation. Even though, among these influence function-based methods, our IP Ensemble demonstrates competitive efficacy with the best or the second best on these three datasets. These datasets cover a range of tasks from marketing prediction and face attribute recognition to toxicity detection, providing a robust evaluation of the model’s defense mechanisms against adaptive adversaries.
@@ 8 Conclusion
In this paper, we revisited and simplified the TracIn method into the Inner Product (IP) formulation, which substitutes the inverse of the Hessian matrix with an identity matrix offers a practical and computationally efficient solution to estimating sample influence. Based on that, we extended our IP to measure the sample influence on fairness and robustness. Continually, we enhanced the generalization of IP by introducing an ensemble strategy. We verified the effectiveness of IP on synthetic datasets and extensive evaluations on noisy label detection, data curation for fair NLP model fine-tuning, and defense against adaptive adversarial attacks. Overall, our IP Ensemble highlighted the potential of simple, yet powerful, approximations in influence estimation for practical use.
Acknowledgment
We gratefully acknowledge the support of the Google TPU Builders Program, which provided support and access to computational resources, and Google Tunix library for post-training functionality and flexibility to enable this work. We also thank the anonymous reviewers and the area chair for their valuable feedback and constructive suggestions.
@@ Acknowledgment
Appendix A Additional Experimental Results and Analysis
A.1 IP for Measuring Fairness and Robustness